Islamabad visited by 53 students, Lahore visited by 52 students, Murree visited by 65 students. Islamabad & Lahore visited by 19 students. Murree & Lahore visted by 28 studerts. Islamabad & Murree visited by 31

students. 38 of them visited exact'y one of the three cities like slamabad, Lohore & Murree. (15)

How many students visited only Islamabad?

How many students visited only Lahore?

How many stuadents visited only Murree?

How many students visited only Islamabad & Lahore but not Murree?

How many students Visited only Muree & Lahore but not Islamabad

How many students visited only Islamabad & Murree but not Lahore?

How many students visited all of the three cities?



1
Expert's answer
2021-01-27T03:47:55-0500

Imagine the number of people who visited these cities in the form of Euler circles:


"X_1+X_5+X_2+X_3=53"

"X_5+X_4+X_3+X_6=52"

"X_2+X_3+X_6+X_7=65"

"X_5+X_3=19"

"X_3+X_6=28"

"X_2+X_3=31"

"X_1+X_4+X_7=38" , therefore

"X_4+X_6=33"

"X_2+X_7=37"

"X_1+X_5=22" , therefore

"X_1+X_5+X_2+X_7+X_4+X_6-" "(X_1+X_4+X_7)=33+37+22-38"

"X_5+X_6+X_2=54"

We summarize the first three equations:

"X_1+X_5+X_2+X_3+ X_5" "+X_4+X_3+X_6+ X_2+X_3+X_6+X_7" "=53+52+65"

"(X_1+X_4+X_7)" "+( 2\times X_5+2\times X_6+2\times X_2)" "+3\times X_3=170"

"3\times X_3=170-108-38"

"X_3=8" - students visited all three cities, therefore

"X_5=11" -students visited only Islamabad & Lahore

"X_6=20" -students Visited only Murree & Lahore

"X_2=23" - students visited only Islamabad & Murree

"X_1=34-23=9" - students visited only Islamabad

"X_4=33-20=13" - students visited only Lahore

"X_7=34-20=14" - students visited only Murree